Consult Multiple-Concept Example 5 in preparation for this problem. The velocity of a diver just before hitting the water is -11.9 m/s, where the minus sign indicates that her motion is directly downward. What is her displacement during the last 1.23 s of the dive?
Answer #1

find the diver's velocity 1.23s before she hits the water.

We know:
v = -11.9 m/s
t = 1.23 s
v0 = ?
g = 9.8 m/s^2

So:
v = v0 - gt
-11.9 = v0 - (9.8)(1.23)
v0 = 0.154 m/s

Now we can solve for displacement with the equation:

V^2 = V0^2 - 2gd, where d = displacement, so:
(-11.9)^2 = (0.154)^2 - 2(9.8)(d)
d = -7.224 m = displacement during last 1.23s of dive